Session: Oct 2009 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Shanghai (Tao An at Fri Oct 23 12:09:23 2009) Success :

Medicina (Sergio Mariotti at Fri Oct 23 16:31:14 2009) Success :
no known problems. Weather: rain, wind

Westerbork (Geert Kuper at Sat Oct 24 08:31:58 2009) Success :
Wb used in phased array mode. No known problems

Onsala (Michael Lindqvist at Sat Oct 24 10:20:03 2009) Success :
No known problems.

Yebes (Susana Garcia at Sun Oct 25 09:11:09 2009) Success :
Not known problems.

Torun (Kaz Borkowski at Mon Oct 26 08:06:08 2009) Success :
No known problems.

Effelsberg (Uwe Bach at Tue Oct 27 08:22:09 2009) Minor failures : ANT
Harddrive failure of the antenna control computer caused a delay of 3 hours. Started observing at 11:30 UT.

Noto (P. Cassaro at Fri Oct 30 12:57:18 2009) Failure : RECO
The DP was erased with a SDK version and used with another. In this way the DP ia seen as full. The experiment has finished before we were able to find out a solution.

Jodrell_Bank (Alastair Gunn at Wed Nov 18 14:07:34 2009) Success :
The Lovell Telescope was used (Jb1). No known problems.

Cambridge (Alastair Gunn at Wed Nov 18 14:08:13 2009) Success :
Darnhall and Knockin telescopes were used. No known problems.